A coin is tossed 10 times. Find the probability that at least three are tails.



1
Expert's answer
2022-05-05T17:06:25-0400

Let "X=" the number of tails: "X\sim Bin (n, p)."

Given "p=0.5, q=1-p=0.5,n=10."


"P(X\ge 3)=1-P(X=0)-P(X=1)"

"-P(X=2)=1-\dbinom{10}{0}(0.5)^0(0.5)^{10-0}"

"-\dbinom{10}{1}(0.5)^1(0.5)^{10-1}-\dbinom{10}{2}(0.5)^2(0.5)^{10-2}"

"=0.9453125"